• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gilman

    Log entry time 22:12:35 on November 4,2008

    Entry number 246741

    keyword=moved spectrometer

    As best as we could determine, the spectrometer had been
    at 15.999 degrees, not 16.003 - the vernier had been poorly read and put into the gui.

    We are now at a floor mark of 46.5 and a vernier of ~+7.8mm, corresponding to an angle of 46.366 degrees.